A psychologist believes that it will take at least an hour for certain disturbed 

children to learn a task. A random sample of 30 of these children results in a 

mean of 50 minutes to learn the task. Should the psychologist modify her belief 

at the 0.01 level if the population standard deviation can be assumed to be 15

minutes?


1
Expert's answer
2021-05-04T12:44:19-0400

M = 50

Z = 2.58

sM = "\sqrt{(15^2/30)}" = 2.74


μ = M ± Z(sM)

μ = 50 ± 2.58"\times"2.74

μ = 50 ± 7.05 (Lower limit is 50 - 7.05 = 42.95, which is more than 30 minutes)


Answer: the belief that children need more than 30 minutes should not be modified, as they need at least 42.95 minutes.